# corresponding author

### v1

@@ 1 Introduction
Artificial intelligence (AI) has been widely adopted in chatbot-based advisory systems. In Vietnam, Phuoc et al. [4] employed Rasa to build a regulation management chatbot capable of intent detection, entity extraction, and dialogue management; while effective for FAQs, its heavy reliance on a small dataset limited robustness for unseen queries. Similarly, Luong and Luong [14] developed an NLP/ML chatbot for student services, providing fast responses but struggling with policy-specific reasoning. Recent advances in large language models (LLMs) [1, 18, 22] such as ChatGPT and Gemini overcome these issues by modeling complex semantics and sustaining multi-turn dialogue, offering valuable insights for academic advisory systems. Complementary techniques have also emerged: Retrieval-Augmented Generation (RAG) improves factual grounding but falters with ambiguous queries, whereas Knowledge Graphs (KGs) capture structured policy relations but lack generative flexibility. To address these gaps, we integrate both in GraphRAG, achieving precise, context-aware, and explainable responses to complex regulation-related questions.
We present REBot, an LLM-powered chatbot that offers CTU students [15] fast, accurate, and up-to-date answers on academic regulations. At its core is CatRAG, a hybrid framework combining RAG for factual reliability with GraphRAG for structured contextual reasoning. To meet the dual need for direct responses and explanatory context, CatRAG integrates NER for semantic enrichment and a routing classifier for precise subgraph selection, enhancing both accuracy and efficiency. Our contributions include: (1) a curated CTU Academic Regulation Dataset optimized for retrieval; (2) a three-tier Knowledge Graph linking categories, chunks, and entities for explainable reasoning; (3) an intent classifier for query routing; (4) CatRAG, a category-guided extension of RAG with new Construct and Query algorithms; and (5) the REBot web application, enabling real-time access to regulation guidance.
The primary goal of GraphRAG is not simply retrieving top-ranked answers, as in conventional RAG, but supplying surrounding contextual knowledge that enriches responses and yields more comprehensive advice. Yet, applied in isolation, GraphRAG often underperforms standard RAG [7]. Our approach therefore seeks both to generate accurate answers and to integrate relevant context, enhancing the overall quality of academic advising.
@@ 2.2 Query-Oriented Text Classification for Academic Regulations
Classifying student queries into thematic categories is essential for improving access to academic regulations. While traditional methods (e.g., Naïve Bayes, KNN, SVM) are widely used [17, 13], they struggle with the linguistic complexity of natural language. In this work, we take advantage of deep learning models (e.g., fasttext) for the classification model. Formally, given queries $\mathcal{X}$ and categories $\mathcal{Y}=\{y_{1},\dots,y_{k}\}$, a classifier $\Phi:\mathcal{X}\to\mathcal{Y}$ maps each query to its most relevant category, enabling intent inference and accurate responses. To ensure coverage, regulatory data are extracted from dispersed PDFs and continuously updated from CTU’s official sources.
@@ 2.3 Semantic Enrichment with NER
Named Entity Recognition (NER) is a central task in information extraction, identifying entities such as people, organizations, and locations [6, 2]. A typical NER pipeline involves segmentation, tokenization, POS tagging, and entity detection, transforming unstructured text into structured representations [11]. In this study, we highlight POS tagging as it provides syntactic cues that constrain entity search, reduce false positives, and improve accuracy [3, 16]. For Vietnamese, we employ Underthesea [^2] https://github.com/undertheseanlp/underthesea, a robust deep learning toolkit with efficient POS tagging [20]. Formally, given a token sequence $\mathit{T}=(t_{1},\dots,t_{n})$, the NER function maps $\mathbf{e}=\mathcal{N}(\mathit{T})=(e_{1},\dots,e_{n}),\quad e_{i}\in\mathcal{E},$ where $\mathcal{E}$ is the entity label set. This formulation underscores POS tagging as a vital step that enhances NER efficiency and accuracy, thereby improving chatbot responses. The integration of RAG and GraphRAG as the framework’s core will be detailed in the following sections.
@@ 2.4 Baseline/General Responses using RAG
Large language models (LLMs), while fluent and generalizable, are prone to AI hallucination [^3] https://cloud.google.com/discover/what-are-ai-hallucinations. Retrieval-augmented methods mitigate this by grounding generation in external knowledge such as corpora, PDFs, or databases. Formally, given a corpus $\mathcal{D}=\{D_{1},\dots,D_{n}\}$ segmented into chunks $\mathcal{C}=\{c_{1},\dots,c_{m}\}$, an embedding function $f_{\mathrm{emb}}$ maps each chunk into $\mathbb{R}^{d}$, producing a vector store $\mathcal{V}=\{(c_{j},f_{\mathrm{emb}}(c_{j}))\mid c_{j}\in\mathcal{C}\},$ which supports efficient similarity-based retrieval. We adopt RAG as the primary mechanism, as it anchors responses in semantically relevant evidence, reducing hallucination and improving factual reliability [12]. This retrieval–generation paradigm provides the foundation for GraphRAG, which further extends capability with structured graph knowledge.
@@ 2.5 Enhanced/Extended Responses using GraphRAG
GraphRAG [8] extends RAG by integrating structured knowledge from knowledge graphs (KGs) into retrieval and generation. Unlike standard RAG, which relies on unstructured corpus, GraphRAG enables entity-centric reasoning and multi-hop traversal over relational structures [10, 12]. As an overview, Figure 1 illustrates the structure of our knowledge graph; its formal definition and role within CatRAG will be detailed in the algorithm section. This structure supports fact retrieval, symbolic reasoning, and semantic search [9, 23, 21], thereby enhancing factual consistency and verifiability [24, 5]. Hence, we adopt GraphRAG for extended responses for it enriches search with structured, entity-level knowledge, allowing deeper reasoning beyond unstructured retrieval alone.
@@ 3 REBot Framework
The framework comprises five interrelated components: (1) Knowledge Extraction, where PDFs are processed with Docling [^4] https://github.com/docling-project/docling for accurate text parsing, cleaned, segmented, and enriched via NER and relation extraction with Underthesea [^5] https://github.com/undertheseanlp/underthesea, then embedded using PhoBERTv2 [19] for reliable Vietnamese representations; (2) Student Affairs Classification ($\Phi$), in which queries are categorized into five regulation domains using fastText [25], chosen for its efficiency and accuracy on short Vietnamese texts; (3) REBot Knowledge Graph, a Neo4j vector-enabled database organizing knowledge into five domain graphs with three layers (categories, contexts, entities), capturing complex student–affairs relations; (4) Vectorization and Embeddings, where LlamaIndex [^6] https://www.llamaindex.ai/ builds a vector index for RAG, combined with PhoBERTv2 embeddings for semantic similarity, and GraphRAG integration to unify vector retrieval with graph structures; and (5) Response Refinement, where Mistral [^7] https://console.mistral.ai/api-keys and ChatGPT [^8] https://platform.openai.com/docs/api-reference refine retrieved content into fluent, context-aware answers.
@@ 10    $label_{c}\leftarrow\Phi(c_{\text{text}})$
11     attach $c$ to category $label_{c}$ in $G$
@@ 5.1 Dataset and Implementation Environment
We compiled two distinct datasets to support our study. The first dataset comprises 1,319 question–answer (Q&A) pairs, carefully curated to evaluate the chatbot’s ability to address inquiries related to academic regulations. The second dataset contains 3,256 questions developed for training a FastText-based classification model. This classification dataset was divided into training and testing subsets using an 80:20 split. Most of the data were sourced from the official departmental websites of Can Tho University [^9] https://www.ctu.edu.vn/don-vi-truc-thuoc.html and other reputable sources. To ensure reliability, all datasets underwent rigorous manual verification. Furthermore, all responses were authored by academic support experts at Can Tho University. The training data and source code of our implementation are publicly available on GitHub [^10] https://github.com/tamB2203579/GraphRAG. The experiments were conducted on a computer equipped with an Intel(R) Core (TM) i5-12500H 4.5 GHz processor, 16 GB of RAM, 6 GB of vRAM and running Windows 11 OS. Next, we will present the experimental results of our approach in the subsequent subsection.

### v2

# REBot: From RAG to CatRAG with Semantic Enrichment and Graph Routing

@@ 1 Introduction
Artificial intelligence (AI) has been widely adopted in chatbot-based advisory systems. In Vietnam, Phuoc et al. [hueuniversity] employed Rasa to build a regulation management chatbot capable of intent detection, entity extraction, and dialogue management; while effective for FAQs, its heavy reliance on a small dataset limited robustness for unseen queries. Similarly, Luong and Luong [luong2025chatbot] developed an NLP/ML chatbot for student services, providing fast responses but struggling with policy-specific reasoning. Recent advances in large language models (LLMs) [YAO2024100211, NAZIR2023100022, inproceedings] such as ChatGPT and Gemini overcome these issues by modeling complex semantics and sustaining multi-turn dialogue, offering valuable insights for academic advisory systems. Complementary techniques have also emerged: Retrieval-Augmented Generation (RAG) improves factual grounding but falters with ambiguous queries, whereas Knowledge Graphs (KGs) capture structured policy relations but lack generative flexibility. To address these gaps, we integrate both in GraphRAG, achieving precise, context-aware, and explainable responses to complex regulation-related questions.
We present REBot, an LLM-powered chatbot that offers CTU students [ma2024racos] fast, accurate, and up-to-date answers on academic regulations. At its core is CatRAG, a hybrid framework combining RAG for factual reliability with GraphRAG for structured contextual reasoning. To meet the dual need for direct responses and explanatory context, CatRAG integrates NER for semantic enrichment and a routing classifier for precise subgraph selection, enhancing both accuracy and efficiency. Our contributions include: (1) a curated CTU Academic Regulation Dataset optimized for retrieval; (2) a three-tier Knowledge Graph linking categories, chunks, and entities for explainable reasoning; (3) an intent classifier for query routing; (4) CatRAG, a category-guided extension of RAG with new Construct and Query algorithms; and (5) the REBot web application, enabling real-time access to regulation guidance.
The primary goal of GraphRAG is not simply retrieving top-ranked answers, as in conventional RAG, but supplying surrounding contextual knowledge that enriches responses and yields more comprehensive advice. Yet, applied in isolation, GraphRAG often underperforms standard RAG [han2025rag]. Our approach therefore seeks both to generate accurate answers and to integrate relevant context, enhancing the overall quality of academic advising.
@@ 2.2 Query-Oriented Text Classification for Academic Regulations
Classifying student queries into thematic categories is essential for improving access to academic regulations. While traditional methods (e.g., Naïve Bayes, KNN, SVM) are widely used [mironczuk2018recentrec, li2022survey], they struggle with the linguistic complexity of natural language. In this work, we take advantage of deep learning models (e.g., fasttext) for the classification model. Formally, given queries $\mathcal{X}$ and categories $\mathcal{Y}=\{y_{1},\dots,y_{k}\}$, a classifier $\Phi:\mathcal{X}\to\mathcal{Y}$ maps each query to its most relevant category, enabling intent inference and accurate responses. To ensure coverage, regulatory data are extracted from dispersed PDFs and continuously updated from CTU’s official sources.
@@ 2.3 Semantic Enrichment with NER
Named Entity Recognition (NER) is a central task in information extraction, identifying entities such as people, organizations, and locations [goyal2018recent, 8999622]. A typical NER pipeline involves segmentation, tokenization, POS tagging, and entity detection, transforming unstructured text into structured representations [huyen2016vlsp]. In this study, we highlight POS tagging as it provides syntactic cues that constrain entity search, reduce false positives, and improve accuracy [chiche2022part, minh2018featurerich]. For Vietnamese, we employ Underthesea [^2] https://github.com/undertheseanlp/underthesea, a robust deep learning toolkit with efficient POS tagging [nguyen2017word]. Formally, given a token sequence $\mathit{T}=(t_{1},\dots,t_{n})$, the NER function maps $\mathbf{e}=\mathcal{N}(\mathit{T})=(e_{1},\dots,e_{n}),\quad e_{i}\in\mathcal{E},$ where $\mathcal{E}$ is the entity label set. This formulation underscores POS tagging as a vital step that enhances NER efficiency and accuracy, thereby improving chatbot responses. The integration of RAG and GraphRAG as the framework’s core will be detailed in the following sections.
@@ 2.4 Baseline/General Responses using RAG
Large language models (LLMs), while fluent and generalizable, are prone to AI hallucination [^3] https://cloud.google.com/discover/what-are-ai-hallucinations. Retrieval-augmented methods mitigate this by grounding generation in external knowledge such as corpora, PDFs, or databases. Formally, given a corpus $\mathcal{D}=\{D_{1},\dots,D_{n}\}$ segmented into chunks $\mathcal{C}=\{c_{1},\dots,c_{m}\}$, an embedding function $f_{\mathrm{emb}}$ maps each chunk into $\mathbb{R}^{d}$, producing a vector store $\mathcal{V}=\{(c_{j},f_{\mathrm{emb}}(c_{j}))\mid c_{j}\in\mathcal{C}\},$ which supports efficient similarity-based retrieval. We adopt RAG as the primary mechanism, as it anchors responses in semantically relevant evidence, reducing hallucination and improving factual reliability [lewis2021retrievalaugmentedgenerationknowledgeintensivenlp]. This retrieval–generation paradigm provides the foundation for GraphRAG, which further extends capability with structured graph knowledge.
@@ 2.5 Enhanced/Extended Responses using GraphRAG
GraphRAG [han2024retrieval] extends RAG by integrating structured knowledge from knowledge graphs (KGs) into retrieval and generation. Unlike standard RAG, which relies on unstructured corpus, GraphRAG enables entity-centric reasoning and multi-hop traversal over relational structures [hu2024grag, lewis2021retrievalaugmentedgenerationknowledgeintensivenlp]. As an overview, Figure 1 illustrates the structure of our knowledge graph; its formal definition and role within CatRAG will be detailed in the algorithm section. This structure supports fact retrieval, symbolic reasoning, and semantic search [Hogan_2021, peng2023knowledge, pan2024unifying], thereby enhancing factual consistency and verifiability [sun2023think, fatemi2023talk]. Hence, we adopt GraphRAG for extended responses for it enriches search with structured, entity-level knowledge, allowing deeper reasoning beyond unstructured retrieval alone.
@@ 3 REBot Framework
The framework comprises five interrelated components: (1) Knowledge Extraction, where PDFs are processed with Docling [^4] https://github.com/docling-project/docling for accurate text parsing, cleaned, segmented, and enriched via NER and relation extraction with Underthesea [^5] https://github.com/undertheseanlp/underthesea, then embedded using PhoBERTv2 [nguyen2020phobert] for reliable Vietnamese representations; (2) Student Affairs Classification ($\Phi$), in which queries are categorized into five regulation domains using fastText [Yao2020TextCM], chosen for its efficiency and accuracy on short Vietnamese texts; (3) REBot Knowledge Graph, a Neo4j vector-enabled database organizing knowledge into five domain graphs with three layers (categories, contexts, entities), capturing complex student–affairs relations; (4) Vectorization and Embeddings, where LlamaIndex [^6] https://www.llamaindex.ai/ builds a vector index for RAG, combined with PhoBERTv2 embeddings for semantic similarity, and GraphRAG integration to unify vector retrieval with graph structures; and (5) Response Refinement, where Mistral [^7] https://console.mistral.ai/api-keys and ChatGPT [^8] https://platform.openai.com/docs/api-reference refine retrieved content into fluent, context-aware answers.
@@ 10    $label_{c}\leftarrow\Phi(c_{\text{text}})$
11     attach $c$ and $v_{c}$ to category $label_{c}$ in $G$
@@ 5.1 Dataset and Implementation Environment
We compiled two distinct datasets to support our study. The first dataset comprises 1,319 question–answer (Q&A) pairs, carefully curated to evaluate the chatbot’s ability to address inquiries related to academic regulations. The second dataset contains 3,256 questions developed for training a FastText-based classification model. This classification dataset was divided into training and testing subsets using an 80:20 split. Most of the data were sourced from the official departmental websites of Can Tho University [^9] https://www.ctu.edu.vn/don-vi-truc-thuoc.html and other reputable sources. To ensure reliability, all datasets underwent rigorous manual verification. Furthermore, all responses were authored by academic support experts at Can Tho University. The training data and source code of our implementation are publicly available on GitHub [^10] https://github.com/tamB2203579/CatRAG. The experiments were conducted on a computer equipped with an Intel(R) Core (TM) i5-12500H 4.5 GHz processor, 16 GB of RAM, 6 GB of vRAM and running Windows 11 OS. Next, we will present the experimental results of our approach in the subsequent subsection.